# TLUE: A Tibetan Language Understanding Evaluation Benchmark

### v3

@@ 5.1 Performance on Ti-MMLU
As shown in Table 1(a), most LLMs perform below the random baseline (25%), underscoring the challenge of Tibetan language understanding. Claude-3.5-Sonnet Anthropic (2024) achieves the highest accuracy, surpassing the baseline by 10.6 percentage points in CAA. Among proprietary LLMs, the GPT Achiam et al. (2023); OpenAI (2024) series underperforms, whereas open-source models like DeepSeek-V3 DeepSeek-AI et al. (2024) exceed the baseline. For reasoning-optimized LLMs, DeepSeek-R1 DeepSeek-AI et al. (2025) surpasses random performance in CAA but underperforms in DA, highlighting Tibetan’s difficulty as a low-resource language.
@@
STEM remains the most challenging category, whereas most models perform best in Social Sciences. However, LlaMA-3.1-405B Dubey et al. (2024) excels in Other, and Qwen-2.5-72b Team (2024) in Humanities, indicating that structured reasoning tasks in Tibetan pose challenges for LLMs, while general knowledge and socially contextualized tasks are relatively easier.
The detial data is released in Appendix Table 15 and Appendix Table 16.
@@ 5.2 Performance on Ti-SafetyBench
As shown in Table 1(b), most LLMs perform below the random baseline (36.7%), indicating significant challenges in aligning models with safety principles in Tibetan. Claude-3.5-Sonnet Anthropic (2024) achieves the highest accuracy (58.5% in CAA), whereas GPT-4O Achiam et al. (2023) and GPT-3.5-Turbo Achiam et al. (2023) underperform, with the latter significantly below random.
Among open-source models, DeepSeek-V3 DeepSeek-AI et al. (2024) surpasses the baseline in both DA and CAA, demonstrating strong safety alignment. LlaMA-3.1-405B Dubey et al. (2024) approaches the baseline, while smaller models like LlaMA-3.1-8B Dubey et al. (2024) and Qwen-2.5-72B Team (2024) perform poorly. Reasoning-optimized models such as DeepSeek-R1 DeepSeek-AI et al. (2025) excel in CAA but struggle in DA, whereas O1-mini OpenAI (2024) consistently underperforms.
@@ 5.3.1 Performance Degradation
All LLMs experience substantial accuracy drops from CMMLU Li et al. (2023) to Ti-MMLU. Qwen-2.5-72B Team (2024) declines from 84.70% $\rightarrow$ 16.50%, GPT-4 Achiam et al. (2023) from 68.90% $\rightarrow$ 17.51%, and ChatGPT Achiam et al. (2023); OpenAI (2024) from 53.22% $\rightarrow$ 3.42%, falling below the random baseline (25%), underscoring the difficulty of adapting LLMs to low-resource languages.
@@ 5.3.2 Domain Performance Shifts
While accuracy declines across all domains, the highest-performing categories differ between languages. In CMMLU Li et al. (2023), LLMs excel in "Other" or "China-specific" categories, whereas in Ti-MMLU, "Social Sciences" and "Humanities" are strongest. For instance, Qwen-2.5-72B Team (2024) achieves the highest accuracy in "Other" (87.35%) on CMMLU Li et al. (2023) but shifts to "Humanities" (17.88%) on Ti-MMLU. LlaMA-3.1-70B Dubey et al. (2024) transitions from "Other" (74.72%) in CMMLU Li et al. (2023) to "Social Sciences" (26.31%) in Ti-MMLU. GPT-4’s Achiam et al. (2023) best-performing category changes from "Other" (73.16%) in CMMLU Li et al. (2023) to "Social Sciences" (19.69%) in Ti-MMLU (GPT-4O Achiam et al. (2023)). These shifts suggest that models retain general knowledge better but struggle with structured reasoning in Tibetan.
@@ 5.3.4 Ranking Inconsistencies
LLM rankings in CMMLU Li et al. (2023) do not consistently translate to Ti-MMLU. For instance, LlaMA-3.1-70B Dubey et al. (2024) ranks below Qwen-2.5-72B Team (2024) in CMMLU Li et al. (2023) but surpasses it in Ti-MMLU, indicating that strong performance in high-resource languages does not necessarily predict effectiveness in low-resource settings.
@@ 5.4 Safety in Low-Resource Languages
(a) Performance Comparison of Qwen-2.5-72B Team (2024) from CMMLU Li et al. (2023) to Ti-MMLU
@@
(c) Performance Comparison of Qwen-2.5 (Chat) Team (2024) from SafetyBench Zhang et al. (2023) to Ti-SafetyBench
@@ 5.4.1 Performance Degradation in Tibetan
LLMs show a significant decline in safety alignment on Ti-SafetyBench, with most falling below the random baseline (36.7%). GPT-4 Achiam et al. (2023) achieves 89.2% on SafetyBench Zhang et al. (2023), but GPT-4O Achiam et al. (2023), despite being a more advanced model, drops to 32.9% on Ti-SafetyBench, particularly in Ethical and Moral Reasoning (92.6% $\rightarrow$ 24.8%). GPT-3.5-Turbo Achiam et al. (2023) experiences a sharper decline (80.4% $\rightarrow$ 11.6%), while Qwen-chat-7B Team (2024) decreases from 77.4% to 30.2%. These findings suggest that LLMs struggle with safety alignment in low-resource languages, often underperforming compared to random selection.
@@ 5.4.2 Variability in Safety Categories
Similarly, Qwen-chat-7B Team (2024) outperforms random in EM and IA on SafetyBench Zhang et al. (2023), but Qwen-2.5-7B Team (2024) retains its best performance only in EM on Ti-SafetyBench. Despite updates, both models show significant degradation in Tibetan.
GPT-3.5-Turbo Achiam et al. (2023) falls well below the random baseline on Ti-SafetyBench, while GPT-4O Achiam et al. (2023) and Qwen-2.5-7B Team (2024) exceed or match the baseline in some categories. These results suggest that factors such as enhanced pretraining data, multilingual adaptation, and fine-tuning may contribute to better transferability, but safety alignment in low-resource settings remains a significant challenge.
@@ 5.5.3 Generalization of Reasoning Models in Low-Resource Languages
As shown in Appendix C Table 12, DeepSeek-R1 Ranathunga et al. (2023) consistently outperforms DeepSeek-V3 DeepSeek-AI et al. (2024) in conditional accuracy. Additionally, O1-mini OpenAI (2024) outperforms GPT-4O Achiam et al. (2023) in STEM tasks, suggesting that stronger reasoning capabilities enhance performance in specialized tasks. However, O1-mini OpenAI (2024) underperforms compared to GPT-4O Achiam et al. (2023) in other domains, indicating that while reasoning optimization improves performance in low-resource settings, it is highly dependent on model architecture and optimization strategies. These results highlight the importance of both reasoning capabilities and generalization in low-resource language adaptation.
@@ 5.6.1 Effect: Scale on Tibetan Understanding
LLMs generally yield better accuracy on Ti-MMLU, though improvements are inconsistent. LlaMA-3.1-405B Dubey et al. (2024) slightly outperforms LlaMA-3.1-70B Dubey et al. (2024), while Qwen-2.5-32B Team (2024) surpasses Qwen-2.5-72B Team (2024). Notably, Qwen-2.5-7B Team (2024) achieves comparable or superior accuracy to Qwen-2.5-72B Team (2024) in several domains, suggesting that model scaling alone does not guarantee enhanced performance in low-resource languages.
@@ 5.6.2 Impact: Scale on Safety Alignment
LLMs show improved performance on Ti-SafetyBench, particularly in categories involving complex ethical or factual reasoning. However, Qwen-2.5-72B Team (2024) significantly outperforms Qwen-2.5-32B Team (2024), indicating that scaling benefits safety alignment more than general understanding. In contrast, LlaMA-3.1-70B Dubey et al. (2024) shows only marginal improvement over LlaMA-3.1-8B Dubey et al. (2024), suggesting that model architecture and adaptation strategies are crucial for safety alignment in low-resource settings.
@@ 5.6.3 Challenge: Scaling for Low-Resource
While LLMs generally perform better, their advantage is less pronounced in Tibetan compared to high-resource languages. Smaller LLMs, such as Qwen-2.5-7B Team (2024), achieve competitive results in specific tasks, indicating that effective pretraining and fine-tuning strategies can mitigate the limitations of smaller models in low-resource environments.
@@
Figure 4: Average accuracy on TLUE across different model scales for LlaMA-3.1 Dubey et al. (2024) and Qwen-2.5 Team (2024).
@@ 6 Conclusion
7 Limitation
The limitation of TLUE is the limited coverage of Tibetan cultural and folk knowledge. While the benchmark includes diverse linguistic and domain-specific tasks, it lacks evaluation datasets tailored to Tibetan folklore, traditional customs, and region-specific knowledge. As cultural understanding is crucial for real-world applications, future work should expand TLUE to better assess LLMs’ capabilities in these areas.
@@ Acknowledgments
This work was supported in part by the National Science and Technology Major Project under Grant 2022ZD0116100, in part by the Tibetan Natural Science Foundation under Grant XZ202201ZR0054G.
@@ Appendix A Appendix: Details of TLUE
As shown in Figure 5, TLUE consists of 2 sub-benchmarks: Ti-MMLU and Ti-SafetyBench. Together, these two components offer a comprehensive evaluation of both the general knowledge proficiency and safety alignment of LLMs in Tibetan. TLUE emphasizes challenges unique to low-resource languages and supports zero-shot and few-shot settings to facilitate cross-model comparisons without relying on language-specific tuning.
@@ A.1 Ti-MMLU
As shown in Firgure 6, Ti-MMLU contains 67 subtasks, covering multiple subject areas from middle school to university and even professional examinations, such as mathematics, physics, history, law, medicine, engineering, philosophy, literature, etc., covering the unique local knowledge system in Tibetan areas, such as college entrance examinations, teacher qualification certificates, medical examinations, etc., and is particularly suitable for evaluating the LLM’s mastery of Tibetan language context and professional knowledge.
@@ A.3 Human Evaluation
In the early days, we used Google Translate and Claude-3.5-Sonnet Anthropic (2024) for translation, which was then optimized by our experts. As shown in Table 2, the translation quality is poor, whether it is manually evaluated or using the BLEU matrix Papineni et al. (2002).
@@ Appendix B Appendix: Algorithm of DA & CCA
6:     $valid\_choices\leftarrow valid\_choices\_map[option\_num]$triangleright Set of valid answer options
7:     for each $excluded$ in $exclude\_list$ do
8:         Remove $excluded$ from $model\_result$
@@
11:     if $|found\_choices|=1$ then
12:         return $found\_choices[0]$triangleright Return extracted answer if unique
@@
6:     $valid\_choices\leftarrow valid\_choices\_map[option\_num]$triangleright Set of valid answer options
7:     for each $excluded$ in $exclude\_list$ do
8:         Remove $excluded$ from $model\_result$
@@
14:     for each $ch$ in $all\_choices$ do
15:         Append $ch$ to $buffer$
16:         if $|buffer|=option\_num$ and $set(buffer)=valid\_choices$ then
17:              Clear $buffer$triangleright Remove one full valid choice combination
18:         else if $|buffer|>option\_num$ then
19:              Append first elements of $buffer$ to $filtered\_choices$
20:              Keep only last $(option\_num-1)$ elements in $buffer$
@@
23:     Append remaining elements in $buffer$ to $filtered\_choices$
@@
26:     if $|unique\_choices|=1$ then
27:         return $list(unique\_choices)$triangleright Return if exactly one unique choice remains
@@
6:     $valid\_choices\leftarrow valid\_choices\_map[option\_num]$triangleright Set of valid answer options $\triangleright$ Filter model-specific reasoning content
7:     if $model\_name$ contains “o1” then
8:         Remove text matching ‘‘Reasoning.*?Reasoned .*? seconds’’ from $model\_result$
9:     else if $model\_name$ contains “DeepSeek-r1” then
10:         Remove text enclosed within ‘‘<think>...</think>’’ from $model\_result$
@@
12:     for each $excluded$ in $exclude\_list$ do
13:         Remove $excluded$ from $model\_result$
@@
16:     if $|found\_choices|=1$ then
17:         return $found\_choices[0]$triangleright Return extracted answer if unique
@@
6:     $valid\_choices\leftarrow valid\_choices\_map[option\_num]$triangleright Set of valid answer options $\triangleright$ Apply model-specific reasoning content filtering
7:     if $model\_name$ contains “o1” then
8:         Remove text matching ‘‘Reasoning.*?Reasoned .*? seconds’’ from $model\_result$
9:     else if $model\_name$ contains “DeepSeek-r1” then
10:         Remove text enclosed within ‘‘<think>...</think>’’ from $model\_result$
@@
12:     for each $excluded$ in $exclude\_list$ do
13:         Remove $excluded$ from $model\_result$
@@
19:     for each $ch$ in $all\_choices$ do
20:         Append $ch$ to $buffer$
21:         if $|buffer|=option\_num$ and $set(buffer)=valid\_choices$ then
22:              Clear $buffer$triangleright Remove one full valid choice combination
23:         else if $|buffer|>option\_num$ then
24:              Append first elements of $buffer$ to $filtered\_choices$
25:              Keep only last $(option\_num-1)$ elements in $buffer$
@@
28:     Append remaining elements in $buffer$ to $filtered\_choices$
@@
31:     if $|unique\_choices|=1$ then
32:         return $list(unique\_choices)$triangleright Return if exactly one unique choice remains

### v5

@@ 5.1 Performance on Ti-MMLU
As shown in Table 1, most LLMs perform below the random baseline (25%), underscoring the challenge of Tibetan language understanding. Claude-3.5-Sonnet Anthropic (2024) achieves the highest accuracy, surpassing the baseline by 10.6 percentage points in CAA. Among proprietary LLMs, the GPT Achiam et al. (2023); OpenAI (2024) series underperforms, whereas open-source models like DeepSeek-V3 DeepSeek-AI et al. (2024) exceed the baseline. For reasoning-optimized LLMs, DeepSeek-R1 DeepSeek-AI et al. (2025) surpasses random performance in CAA but underperforms in DA, highlighting Tibetan’s difficulty as a low-resource language.
@@
STEM remains the most challenging category, whereas most models perform best in Social Sciences. However, LlaMA-3.1-405B Dubey et al. (2024) excels in Other, and Qwen-2.5-72b Qwen Team (2024) in Humanities, indicating that structured reasoning tasks in Tibetan pose challenges for LLMs, while general knowledge and socially contextualized tasks are relatively easier.
The detailed results can be found in Appendix Table 15 and Appendix Table 16.
@@ 5.2 Performance on Ti-SafetyBench
As shown in Table 1, most LLMs perform below the random baseline (36.7%), indicating significant challenges in aligning models with safety principles in Tibetan. Claude-3.5-Sonnet Anthropic (2024) achieves the highest accuracy (58.5% in CAA), whereas GPT-4o Achiam et al. (2023) and GPT-3.5-Turbo Achiam et al. (2023) underperform, with the latter significantly below random.
Among open-source models, DeepSeek-V3 DeepSeek-AI et al. (2024) surpasses the baseline in both DA and CAA, demonstrating strong safety alignment. LlaMA-3.1-405B Dubey et al. (2024) approaches the baseline, while smaller models like LlaMA-3.1-8B Dubey et al. (2024) and Qwen-2.5-72B Qwen Team (2024) perform poorly. Reasoning-optimized models such as DeepSeek-R1 DeepSeek-AI et al. (2025) excel in CAA but struggle in DA, whereas O1-mini OpenAI (2024) consistently underperforms.
@@ 5.3.1 Performance Degradation
All LLMs experience substantial accuracy drops from CMMLU Li et al. (2023) to Ti-MMLU. Qwen-2.5-72B Qwen Team (2024) declines from 84.70% $\rightarrow$ 16.50%, GPT-4 Achiam et al. (2023) from 68.90% $\rightarrow$ 17.51%, and ChatGPT Achiam et al. (2023); OpenAI (2024) from 53.22% $\rightarrow$ 3.42%, falling below the random baseline (25%), underscoring the difficulty of adapting LLMs to low-resource languages.
@@ 5.3.2 Domain Performance Shifts
While accuracy declines across all domains, the highest-performing categories differ between languages. In CMMLU Li et al. (2023), LLMs excel in "Other" or "China-specific" categories, whereas in Ti-MMLU, "Social Sciences" and "Humanities" are strongest. For instance, Qwen-2.5-72B Qwen Team (2024) achieves the highest accuracy in "Other" (87.35%) on CMMLU Li et al. (2023) but shifts to "Humanities" (17.88%) on Ti-MMLU. LlaMA-3.1-70B Dubey et al. (2024) transitions from "Other" (74.72%) in CMMLU Li et al. (2023) to "Social Sciences" (26.31%) in Ti-MMLU. GPT-4’s Achiam et al. (2023) best-performing category changes from "Other" (73.16%) in CMMLU Li et al. (2023) to "Social Sciences" (19.69%) in Ti-MMLU (GPT-4o Achiam et al. (2023)). These shifts suggest that models retain general knowledge better but struggle with structured reasoning in Tibetan.
@@ 5.3.4 Ranking Inconsistencies
LLM rankings in CMMLU Li et al. (2023) do not consistently translate to Ti-MMLU. For instance, LlaMA-3.1-70B Dubey et al. (2024) ranks below Qwen-2.5-72B Qwen Team (2024) in CMMLU Li et al. (2023) but surpasses it in Ti-MMLU, indicating that strong performance in high-resource languages does not necessarily predict effectiveness in low-resource settings.
@@ 5.4 Safety in Low-Resource Languages
(a) Performance Comparison of Qwen-2.5-72B Qwen Team (2024) from CMMLU Li et al. (2023) to Ti-MMLU
@@
(c) Performance Comparison of Qwen-2.5 (Chat) Qwen Team (2024) from SafetyBench Zhang et al. (2023) to Ti-SafetyBench
@@ 5.4.1 Performance Degradation in Tibetan
LLMs show a significant decline in safety alignment on Ti-SafetyBench, with most falling below the random baseline (36.7%). GPT-4 Achiam et al. (2023) achieves 89.2% on SafetyBench Zhang et al. (2023), but GPT-4o Achiam et al. (2023), despite being a more advanced model, drops to 32.9% on Ti-SafetyBench, particularly in Ethical and Moral Reasoning (92.6% $\rightarrow$ 24.8%). GPT-3.5-Turbo Achiam et al. (2023) experiences a sharper decline (80.4% $\rightarrow$ 11.6%), while Qwen-chat-7B Qwen Team (2024) decreases from 77.4% to 30.2%. These findings suggest that LLMs struggle with safety alignment in low-resource languages, often underperforming compared to random selection.
@@ 5.4.2 Variability in Safety Categories
Similarly, Qwen-chat-7B Qwen Team (2024) outperforms random in EM and IA on SafetyBench Zhang et al. (2023), but Qwen-2.5-7B Qwen Team (2024) retains its best performance only in EM on Ti-SafetyBench. Despite updates, both models show significant degradation in Tibetan.
GPT-3.5-Turbo Achiam et al. (2023) falls well below the random baseline on Ti-SafetyBench, while GPT-4o Achiam et al. (2023) and Qwen-2.5-7B Qwen Team (2024) exceed or match the baseline in some categories. These results suggest that factors such as enhanced pretraining data, multilingual adaptation, and fine-tuning may contribute to better transferability, but safety alignment in low-resource settings remains a significant challenge.
@@ 5.5.3 Generalization of Reasoning Models in Low-Resource Languages
As shown in Appendix C Table 12, DeepSeek-R1 DeepSeek-AI et al. (2025) consistently outperforms DeepSeek-V3 DeepSeek-AI et al. (2024) in conditional accuracy. Additionally, O1-mini OpenAI (2024) outperforms GPT-4o Achiam et al. (2023) in STEM tasks, suggesting that stronger reasoning capabilities enhance performance in specialized tasks. However, O1-mini OpenAI (2024) underperforms compared to GPT-4o Achiam et al. (2023) in other domains, indicating that while reasoning optimization improves performance in low-resource settings, it is highly dependent on model architecture and optimization strategies. These results highlight the importance of both reasoning capabilities and generalization in low-resource language adaptation.
@@ 5.6.1 Effect: Scale on Tibetan Understanding
LLMs generally yield better accuracy on Ti-MMLU, though improvements are inconsistent. LlaMA-3.1-405B Dubey et al. (2024) slightly outperforms LlaMA-3.1-70B Dubey et al. (2024), while Qwen-2.5-32B Qwen Team (2024) surpasses Qwen-2.5-72B Qwen Team (2024). Notably, Qwen-2.5-7B Qwen Team (2024) achieves comparable or superior accuracy to Qwen-2.5-72B Qwen Team (2024) in several domains, suggesting that model scaling alone does not guarantee enhanced performance in low-resource languages.
@@ 5.6.2 Impact: Scale on Safety Alignment
LLMs show improved performance on Ti-SafetyBench, particularly in categories involving complex ethical or factual reasoning. However, Qwen-2.5-72B Qwen Team (2024) significantly outperforms Qwen-2.5-32B Qwen Team (2024), indicating that scaling benefits safety alignment more than general understanding. In contrast, LlaMA-3.1-70B Dubey et al. (2024) shows only marginal improvement over LlaMA-3.1-8B Dubey et al. (2024), suggesting that model architecture and adaptation strategies are crucial for safety alignment in low-resource settings.
@@ 5.6.3 Challenge: Scaling for Low-Resource
While LLMs generally perform better, their advantage is less pronounced in Tibetan compared to high-resource languages. Smaller LLMs, such as Qwen-2.5-7B Qwen Team (2024), achieve competitive results in specific tasks, indicating that effective pretraining and fine-tuning strategies can mitigate the limitations of smaller models in low-resource environments.
@@
Figure 4: Average accuracy on TLUE across different model scales for LlaMA-3.1 Dubey et al. (2024) and Qwen-2.5 Qwen Team (2024).
@@ 6 Conclusion
7 Limitation & Future Work
A limitation of TLUE is its limited coverage of Tibetan cultural and folk knowledge. We will propuse TLUE+ to complete the benchmark content of this part.
The TLUE benchmark involves evaluation-only data in the Tibetan language, curated with expert review to ensure cultural and ethical appropriateness. While the benchmark includes safety-critical topics (e.g., ethics, bias, religion), all content was manually vetted by native speakers and domain experts. We believe the potential risks are minimal and primarily relate to possible model misuse in sensitive domains, rather than any harm caused by the dataset itself.
Ethics Statement
In constructing the TLUE benchmark, we adhered to ethical standards throughout. All data is used solely for evaluation, contains no personal or sensitive information, and was manually reviewed by native Tibetan speakers and domain experts to ensure cultural and linguistic integrity. This work aims to promote fair representation of minority languages like Tibetan and avoid any form of cultural bias or discrimination.
@@ Acknowledgments
This work was supported in part by the National Science and Technology Major Project under Grant 2022ZD0116100, in part by the Sichuan Provincial Major Science and Technology Project under Grant 2024ZDZX0012, in part by the National Natural Science Foundation of China under Grants 62276055, 62406062, 62436006 and 62406257,in part by the Sichuan Science and Technology Program under Grant 2023YFG0288, in part by the Natural Science Foundation of Sichuan Province under Grant 2024NSFSC1476, in part by the Tibetan Natural Science Foundation under GrantXZ202201ZR0054G.
@@ Appendix A Appendix: Details of TLUE
As shown in Figure 5, TLUE [^1] License: CC-BY-NC-SA 4.0 consists of 2 sub-benchmarks: Ti-MMLU and Ti-SafetyBench. Together, these two components offer a comprehensive evaluation of both the general knowledge proficiency and safety alignment of LLMs in Tibetan. TLUE emphasizes challenges unique to low-resource languages and supports zero-shot and few-shot settings to facilitate cross-model comparisons without relying on language-specific tuning.
@@ A.1 Ti-MMLU
As shown in Figure 6, Ti-MMLU contains 67 subtasks, covering multiple subject areas from middle school to university and even professional examinations, such as mathematics, physics, history, law, medicine, engineering, philosophy, literature, etc., covering the unique local knowledge system in Tibetan areas, such as college entrance examinations, teacher qualification certificates, medical examinations, etc., and is particularly suitable for evaluating the LLM’s mastery of Tibetan language context and professional knowledge.
@@ A.3 Human Evaluation
In the early days, we used Google Translate and Claude-3.5-Sonnet Anthropic (2024) for translation, which was then optimized by our experts. As shown in Table 2, the translation quality is poor, whether it is manually evaluated or using the BLEU metric Papineni et al. (2002).
@@ Appendix B Appendix: Algorithm of DA & CCA
6:  $valid\_choices\leftarrow valid\_choices\_map[option\_num]$ $\triangleright$ Set of valid answer options
7:  for each excludedexcluded in exclude_listexclude\_list do
8:   Remove excludedexcluded from model_resultmodel\_result
@@
11:  if |found_choices|=1|found\_choices|=1 then
12:   return found_choices[0]found\_choices[0] $\triangleright$ Return extracted answer if unique
@@
6:  $valid\_choices\leftarrow valid\_choices\_map[option\_num]$ $\triangleright$ Set of valid answer options
7:  for each excludedexcluded in exclude_listexclude\_list do
8:   Remove excludedexcluded from model_resultmodel\_result
@@
14:  for each chch in all_choicesall\_choices do
15:   Append chch to bufferbuffer
16:   if |buffer|=option_num|buffer|=option\_num and set(buffer)=valid_choicesset(buffer)=valid\_choices then
17:     Clear bufferbuffer $\triangleright$ Remove one full valid choice combination
18:   else if |buffer|>option_num|buffer|>option\_num then
19:     Append first elements of bufferbuffer to filtered_choicesfiltered\_choices
20:     Keep only last (option_num1)(option\_num-1) elements in bufferbuffer
@@
23:  Append remaining elements in bufferbuffer to filtered_choicesfiltered\_choices
@@
26:  if |unique_choices|=1|unique\_choices|=1 then
27:   return list(unique_choices)list(unique\_choices) $\triangleright$ Return if exactly one unique choice remains
@@
6:  $valid\_choices\leftarrow valid\_choices\_map[option\_num]$ $\triangleright$ Set of valid answer options $\triangleright$ Filter model-specific reasoning content
7:  if model_namemodel\_name contains “o1” then
8:   Remove text matching ‘‘Reasoning.*?Reasoned .*? seconds’’ from model_resultmodel\_result
9:  else if model_namemodel\_name contains “DeepSeek-r1” then
10:   Remove text enclosed within ‘‘<think>...</think>’’ from model_resultmodel\_result
@@
12:  for each excludedexcluded in exclude_listexclude\_list do
13:   Remove excludedexcluded from model_resultmodel\_result
@@
16:  if |found_choices|=1|found\_choices|=1 then
17:   return found_choices[0]found\_choices[0] $\triangleright$ Return extracted answer if unique
@@
6:  $valid\_choices\leftarrow valid\_choices\_map[option\_num]$ $\triangleright$ Set of valid answer options $\triangleright$ Apply model-specific reasoning content filtering
7:  if model_namemodel\_name contains “o1” then
8:   Remove text matching ‘‘Reasoning.*?Reasoned .*? seconds’’ from model_resultmodel\_result
9:  else if model_namemodel\_name contains “DeepSeek-r1” then
10:   Remove text enclosed within ‘‘<think>...</think>’’ from model_resultmodel\_result
@@
12:  for each excludedexcluded in exclude_listexclude\_list do
13:   Remove excludedexcluded from model_resultmodel\_result
@@
19:  for each chch in all_choicesall\_choices do
20:   Append chch to bufferbuffer
21:   if |buffer|=option_num|buffer|=option\_num and set(buffer)=valid_choicesset(buffer)=valid\_choices then
22:     Clear bufferbuffer $\triangleright$ Remove one full valid choice combination
23:   else if |buffer|>option_num|buffer|>option\_num then
24:     Append first elements of bufferbuffer to filtered_choicesfiltered\_choices
25:     Keep only last (option_num1)(option\_num-1) elements in bufferbuffer
@@
28:  Append remaining elements in bufferbuffer to filtered_choicesfiltered\_choices
@@
31:  if |unique_choices|=1|unique\_choices|=1 then
32:   return list(unique_choices)list(unique\_choices) $\triangleright$ Return if exactly one unique choice remains